OpenStudy (anonymous):

a wave moves with speed 300m/s on awire which is under a tention of 500N . find how much the tention must be changed to increase the speed to 312m/s

OpenStudy (anonymous):

v is inversely proportional to the square root of tension. do you need further help?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

since: \(v=\sqrt \frac{T}{m}\)where m is the mass perlength, a constant, rearranging, \(\frac{v_1}{\sqrt T_1}=\frac{v_2}{\sqrt T_2}\) sub it all in.